# Cortex AISQL: A Production SQL Engine for Unstructured Data

### v1

@@ 6.2 Adaptive Model Cascades
Our evaluation methodology executes each query five times to ensure statistical reliability. We measure both execution efficiency (query time) and prediction quality using standard classification metrics: accuracy, precision, recall, and F1 score. All quality assessments use ground-truth labels from the evaluation datasets as reference.
Table 2 reports the mean performance across all datasets, including execution time, speedup relative to the baseline, F1 score with delta () relative to baseline, and precision/recall metrics. All speedup percentages and delta F1 values are computed with respect to the baseline approach.
@@
Performance gains. The cascade approach achieves substantial speedup with mean execution time reduced by 65.5% (from 975.9 to 336.4 seconds), corresponding to a 2.9 speedup. This improvement stems from the cascade’s ability to route the majority of predictions through the lightweight proxy model, reserving the more expensive oracle model only for uncertain cases. For comparison, the proxy model alone executes in 296.2 seconds (3.3 speedup), as expected from its smaller model size, though this comes at a substantial quality cost.
@@
Figure 11 shows per-dataset results, revealing that speedup varies considerably across workloads. The NQ dataset achieves the highest speedup at 5.85, while QUORA and FARL show moderate improvements (1.22 and 1.30 respectively). These variations reflect differences in how the threshold learning algorithm partitions the data – easier workloads where the proxy model exhibits stronger confidence scores yield higher routing rates to the proxy and thus greater speedups.
@@
Quality analysis. The cascade maintains strong prediction quality with only modest mean degradation. F1 score decreases by 4.3% (from 0.812 to 0.777), with balanced precision and recall remaining at 0.784 and 0.794 respectively. This demonstrates that the threshold learning logic effectively identifies rows where the proxy model’s confidence scores are reliable. In contrast, using the proxy model exclusively results in substantially degraded quality, with F1 dropping by 18.8% to 0.659, demonstrating the oracle model’s critical role in handling predictions that fall within the uncertainty region.
@@
As shown in Figure 11, quality impact varies across datasets. NQ maintains near-identical F1 scores (), indicating that most predictions have clear confidence patterns suitable for routing. IMDB and SST2 show moderate degradation ( and respectively), while BOOLQ and QUORA experience larger drops ( and ). These variations suggest that cascade effectiveness depends on dataset characteristics, i.e., tasks where the proxy model’s uncertainty correlates well with prediction difficulty benefit most from the adaptive routing strategy.
@@
Discussion. The cascade offers an attractive balance between the baseline and proxy-only approaches. For applications where near-optimal quality is essential, the baseline provides the highest F1 (0.812) but at 2.9 higher computational cost. For latency-critical workloads where moderate quality is acceptable, the cascade delivers 65.5% speedup while retaining 95.7% of baseline F1 performance. The proxy-only configuration is suitable only when speed is paramount and substantial quality degradation is tolerable.
@@
Beyond the speed-quality trade-off, the cascade system provides two additional control mechanisms. Users can constrain the oracle budget by specifying a maximum number of oracle invocations, ensuring predictable costs regardless of data characteristics. The system also supports precision and recall targets, allowing users to specify desired quality thresholds. However, achieving these targets may require higher delegation rates to the oracle model, reducing the speedup gains.
@@
We also found the cascades algorithm to be resilient to both the degree of parallelism in the SQL execution engine and the number of rows processed by individual workers.
6.3 Query Rewriting for Semantic Joins
@@
We evaluate the AI_CLASSIFY rewrite optimization using eight semantic join benchmarks spanning entity matching (ABTBUY, NASDAQ), document categorization (ARXIV, EURLEX, NYT, CNN), news classification (AG NEWS), and biomedical concept linking (BIODEX). Each benchmark consists of two tables with natural language content, where the join task is to identify semantically related row pairs. For example, in ABTBUY the predicate matches product descriptions from different e-commerce sites that refer to the same item, while in NYT it links news articles to their topical categories. The left table (denoted as in Table 4) and right table () span from 50 to 500 rows, yielding up to 250,000 candidate pairs that must be considered by the naive join approach using filtering.
@@
Table 3 summarizes mean performance across all benchmarks. Table 4 reports the number of LLM calls, total execution time (seconds), and standard classification metrics (precision, recall, F1) for both approaches across datasets of varying cardinalities.
Performance gains. The rewrite achieves dramatic speedups across all benchmarks, ranging from (ARXIV) to (CNN), with a mean speedup of . These improvements stem directly from reducing LLM invocations: the baseline requires calls while the rewrite needs only classifications. For the CNN dataset with 500 rows per table, this reduces 250,000 binary classifications to just 1,000 multi-label classifications, cutting execution time from 4.4 hours to 3.8 minutes. The speedup magnitude correlates strongly with dataset scale—larger Cartesian products yield proportionally greater benefits.
@@
Quality analysis. The impact of the rewrite optimization on prediction quality varies by dataset characteristics. On entity matching tasks with clear semantic signals (ABTBUY), both approaches achieve nearly identical F1 scores ( vs ), demonstrating that the rewrite preserves quality when matches are unambiguous. More remarkably, the rewrite substantially improves quality on datasets where the baseline suffers from poor precision. For NASDAQ, the extremely low precision of the baseline () results in an F1 of , while the rewrite achieves precision and F1. We attribute this to the multi-label classification paradigm: by presenting all candidate labels simultaneously, AI_CLASSIFY enables better comparative reasoning than isolated binary decisions in AI_FILTER. Overall, mean F1 improves by (from to ), with precision increasing from to .
Trade-offs. The rewrite exhibits precision-recall trade-offs that depend on the classification strategy. Datasets like EURLEX and BIODEX show recall degradation ( and , respectively) despite improved precision. This occurs when the model conservatively selects matches from the full label set, prioritizing accuracy over coverage. For applications where recall is critical, hybrid strategies combining both approaches or tuning the classification prompt may be necessary.
7 Related Work
@@
The AI4DB line of research applies ML algorithms to internal database components such as query optimization, indexing, and configuration tuning. Early systems like LEO [stillger2001leo] used feedback learning to refine cost estimates for the DB2 optimizer. Later, OtterTune [vanaken2017ottertune] and CDBTune [zhang2019end] leveraged large-scale supervised and reinforcement learning to automatically tune database parameters, inspiring the vision of self-driving databases [pavlo2017self].
A rich body of work explores learned query processing components. Learned Indexes [kraska2018case] replaced traditional index structures with neural networks, while ALEX [ding2020alex] extended these ideas to support dynamic workloads. In query optimization, learned cardinality estimation models, such as MSCN [kipf2018learned], Naru [yang2019deep], and DeepDB [hilprecht2020deepdb], achieved notable improvements in selectivity prediction by learning correlations across attributes and tables. Reinforcement learning has also been applied to plan search and join ordering, exemplified by ReJOIN [marcus2018deep], Neo [marcus2019neo], and Bao [marcus2021bao]. These works culminate in the SageDB [kraska2019sagedb], which proposes a learned database system where each component is automatically specialized for a given data distribution and workload.
The application of ML to query optimization extends to handling expensive predicates in specialized domains. Systems such as BlazeIt [kang2019blazeit] and TASTI [kang2020tasti] optimize queries involving expensive ML-based predicates using model cascades, semantic indexes, and query-specific proxy models. In particular, model cascades that combine cheap approximate models with expensive accurate ones [kang2020approximate, chen2023frugal, zellinger2025rational] have influenced the design of AISQL’s approach to managing expensive AI operations.
@@
Earlier in-database ML frameworks such as MADLib [hellerstein2012madlib] extended SQL with support for descriptive statistics and simple ML methods. Recent systems go further by integrating semantic and language-model capabilities directly into query processing, for example LOTUS [patel2025semantic] introduces a set of semantic operators for Pandas-like dataframe processing. ThalamusDB [jo2025thalamusdb] explores approximate query processing for multimodal data, supporting natural language predicates over visual, audio, and textual content through zero-shot models combined with relational operators.
Several systems have proposed specialized approaches to LLM-powered SQL processing. SUQL [liu2024suql] augments SQL with answer and summary operators for knowledge-grounded conversational agents, with the focus on row-wise LLM operations for question answering. ZenDB [lin2024towards] optimizes SQL queries for extracting structured data from semi-structured documents with predictable templates. Palimpzest [liu2025palimpzest] provides a declarative framework for LLM-powered data processing with specialized map-like operations and basic cost-based optimizations. UQE [dai2024uqe] studies embedding-based approximations for LLM-powered filters and stratified sampling for aggregations, though it provides best-effort performance without accuracy guarantees.
Beyond SQL-centric approaches, several frameworks have emerged for LLM-powered data processing. Aryn [anderson2024design] offers a Spark-like API with PDF extraction and human-in-the-loop processing capabilities. DocETL [shankar2025docetl] proposes agent-driven pipeline optimization for complex document processing tasks and uses LLM agents to explore task decomposition strategies. EVAPORATE [arora2023language] specializes in extracting semi-structured tables through code synthesis. These systems demonstrate the potential of LLM-powered processing, but often lack formal cost models and provide limited performance guarantees.
@@
8 Conclusions
The proliferation of unstructured data alongside traditional structured datasets has created a fundamental gap in modern data warehouses: users need to perform semantic reasoning over hybrid data but are forced to export data to external systems, orchestrate complex pipelines, and sacrifice the optimization benefits of integrated query processing. This paper presents Snowflake AISQL, a production SQL engine that bridges this gap by integrating LLM capabilities directly into SQL through native semantic operators.
AISQL introduces six semantic operators that compose naturally with traditional relational SQL primitives. These operators enable users to express complex analytical tasks that blend structured queries with semantic reasoning in a single declarative statement, and eliminate the need for external data movement and manual pipeline orchestration.
The core technical challenge we address is making semantic operators efficient at production scale. LLM-based semantic operators are orders of magnitude more expensive than traditional SQL operations, their cost and selectivity are opaque to the query optimizer, and conventional optimization heuristics can produce catastrophically inefficient plans. Our solution comprises three techniques driven by analysis of real customer workloads. AI-aware query optimization treats LLM inference cost as a first-class objective, which leads to efficient operator placement and predicate evaluation order. Adaptive model cascades reduce costs through intelligent routing between lightweight proxy models and powerful oracle models. Query rewriting for semantic joins transforms quadratic-complexity joins into linear multi-label classification task and, as a result, delivers speedups with improved prediction quality.
Our production deployment at Snowflake validates these design choices. Analysis of customer workloads reveals that AI operators consume the dominant share of execution costs and nearly half of queries involve multi-table operations. These observations directly motivate our focus on join optimization and cost-aware planning. Extensive experimental evaluation across diverse benchmarks demonstrates that AISQL successfully delivers both efficiency and quality at production scale.
AISQL represents a step toward unified data processing architectures where semantic reasoning over unstructured content is a first-class database capability rather than an external add-on. By bringing LLM capabilities into the SQL engine with attention to cost optimization, we enable organizations to unlock insights from their data assets through familiar declarative interfaces while maintaining the performance, scalability, and reliability expected from modern data warehouses.

### v2

@@ 6.2. Adaptive Model Cascades
Our evaluation methodology executes each query five times to ensure statistical reliability. We measure both execution efficiency (query time) and prediction quality using standard classification metrics: accuracy, precision, recall, and F1 score. All quality assessments use ground-truth labels from the evaluation datasets as a reference.
Table 2 reports the mean performance across all datasets, including execution time, speedup relative to the baseline, F1 score with delta () relative to the baseline, and precision/recall metrics. All speedup percentages and delta F1 values are computed with respect to the baseline approach.
@@
Performance gains. The cascade approach achieves substantial speedup with mean execution time reduced by 65.5% (from 975.9 to 336.4 seconds), corresponding to a 2.9 speedup. This improvement stems from the cascade’s ability to route the majority of predictions through the lightweight proxy model, reserving the more expensive oracle model only for uncertain cases. For comparison, the proxy model alone runs in 296.2 seconds (3.3 speedup), as expected from its smaller model size, though this comes at a substantial quality cost.
@@
Figure 11 shows per-dataset results, revealing that speedup varies considerably across workloads. The NQ dataset achieves the highest speedup at 5.85, while QUORA and FARL show moderate improvements (1.22 and 1.30 respectively). These variations reflect differences in how the threshold learning algorithm partitions the data. Easier workloads, where the proxy model exhibits stronger confidence scores, yield higher routing rates to the proxy and thus greater speedups.
@@
Quality analysis. The cascade maintains strong prediction quality with only modest mean degradation. The F1 score decreases by 4.3% (from 0.812 to 0.777), with balanced precision and recall remaining at 0.784 and 0.794, respectively. This demonstrates that the threshold learning logic effectively identifies rows where the proxy model’s confidence scores are reliable. In contrast, using the proxy model exclusively results in substantially degraded quality, with F1 dropping by 18.8% to 0.659, demonstrating the critical role of the oracle model in handling predictions that fall within the uncertainty region.
@@
As shown in Figure 11, the impact on quality scores varies across datasets. NQ maintains near-identical F1 scores (), indicating that most predictions have clear confidence patterns suitable for routing. IMDB and SST2 show moderate degradation ( and respectively), while BOOLQ and QUORA experience larger drops ( and ). These variations suggest that cascade effectiveness depends on dataset characteristics, i.e., tasks where the proxy model’s uncertainty correlates well with prediction difficulty benefit most from the adaptive routing strategy.
@@
Discussion. The cascade offers an attractive balance between the baseline and proxy-only approaches. For applications where near-optimal quality is essential, the baseline achieves the highest F1 score (0.812) but at a 2.9 higher computational cost. For latency-critical workloads where moderate quality is acceptable, the cascade delivers 65.5% speedup while retaining 95.7% of baseline F1 performance. The proxy-only configuration is suitable only when speed is paramount and substantial quality degradation is tolerable.
@@
Beyond the speed-quality trade-off, the cascade system provides two additional control mechanisms. Users can restrict the oracle budget by specifying a maximum number of oracle invocations, ensuring predictable costs regardless of data characteristics. The system also supports precision and recall targets, allowing users to specify desired quality thresholds. However, achieving these targets may require higher delegation rates to the oracle model, which reduces the speedup gains.
@@
We also found the cascade algorithm to be resilient to both the degree of parallelism in the SQL execution engine and the number of rows processed by individual workers.
6.3. Query Rewriting for Semantic Joins
@@
We evaluate AI_CLASSIFY rewrite optimization using eight semantic join benchmarks that cover entity matching (ABTBUY, NASDAQ), document categorization (ARXIV, EURLEX, NYT, CNN), news classification (AG NEWS), and biomedical concept linking (BIODEX). Each benchmark consists of two tables with natural language content, where the join task is to identify semantically related row pairs. For example, in ABTBUY the predicate matches product descriptions from different e-commerce sites that refer to the same item, while in NYT it links news articles to their topical categories. The left table (denoted as in Table 4) and the right table () span 50 to 500 rows, yielding up to 250,000 candidate pairs that must be considered by the naive join approach using filtering.
@@
Table 3 summarizes the mean performance across all benchmarks. Table 4 reports the number of LLM calls, the total execution time (seconds), and standard classification metrics (precision, recall, F1) for both approaches across datasets of varying cardinalities.
Table 4 reports the number of LLM calls, the total execution time (seconds), and standard classification metrics (precision, recall, F1) for both approaches across datasets of varying cardinalities. Performance gains. The query rewriting achieves dramatic speedups across all benchmarks, ranging from (ARXIV) to (CNN), with a mean speedup of . These improvements stem directly from reducing LLM invocations: the baseline requires calls, while the rewrite needs only classifications. For the CNN dataset with 500 rows per table, this reduces 250,000 binary classifications to just 1,000 multi-label classifications, cutting execution time from 4.4 hours to 3.8 minutes. The speedup magnitude correlates strongly with dataset scale; larger Cartesian products yield proportionally greater benefits.
@@
Quality analysis. The impact of rewrite optimization on prediction quality varies by dataset characteristics. On entity matching tasks with clear semantic signals (ABTBUY), both approaches achieve nearly identical F1 scores ( vs ), demonstrating that the rewrite preserves quality when matches are unambiguous. More remarkably, the rewrite substantially improves quality on datasets where the baseline suffers from poor precision. For NASDAQ, the extremely low precision of the baseline () results in an F1 score of , while the rewrite achieves a precision of and an F1 score of . We attribute this to the multi-label classification paradigm: by presenting all candidate labels simultaneously, AI_CLASSIFY enables better comparative reasoning than isolated binary decisions in AI_FILTER. The mean F1 improves by (from to ), with precision increasing from to . Overall, the mean F1 improves by (from to ), with precision increasing from to .
Trade-offs. The rewrite exhibits precision-recall trade-offs that depend on the classification strategy. Datasets like EURLEX and BIODEX show recall degradation ( and , respectively) despite improved precision. Recall degradation occurs when the model conservatively selects matches from the full set of labels, prioritizing accuracy over coverage. For applications where recall is critical, hybrid strategies that combine both approaches or tune the classification prompt may be necessary.
7. Related Work
@@
The AI4DB line of research applies ML algorithms to internal database components such as query optimization, indexing, and configuration tuning. Early systems like LEO (stillger2001leo) used feedback learning to refine cost estimates for the DB2 optimizer. Later, OtterTune (vanaken2017ottertune) and CDBTune (zhang2019end) leveraged large-scale supervised and reinforcement learning to automatically tune database parameters, inspiring the vision of self-driving databases (pavlo2017self).
A rich body of work explores learned query processing components. Learned Indexes (kraska2018case) replaced traditional index structures with neural networks, while ALEX (ding2020alex) extended these ideas to support dynamic workloads. In query optimization, learned cardinality estimation models, such as MSCN (kipf2018learned), Naru (yang2019deep), and DeepDB (hilprecht2020deepdb), achieved notable improvements in selectivity prediction by learning correlations across attributes and tables. Reinforcement learning has also been applied to plan search and join ordering, exemplified by ReJOIN (marcus2018deep), Neo (marcus2019neo), and Bao (marcus2021bao). These works culminate in the SageDB (kraska2019sagedb), which proposes a learned database system where each component is automatically specialized for a given data distribution and workload.
The application of ML to query optimization extends to handling expensive predicates in specialized domains. Systems such as BlazeIt (kang2019blazeit) and TASTI (kang2020tasti) optimize queries involving expensive ML-based predicates using model cascades, semantic indexes, and query-specific proxy models. In particular, model cascades that combine cheap approximate models with expensive accurate ones (kang2020approximate; chen2023frugal; zellinger2025rational) have influenced the design of AISQL’s approach to managing expensive AI operations.
@@
Earlier in-database ML frameworks such as MADLib (hellerstein2012madlib) extended SQL with support for descriptive statistics and simple ML methods. Recent systems go further by integrating semantic and language-model capabilities directly into query processing, for example LOTUS (patel2025semantic) introduces a set of semantic operators for Pandas-like dataframe processing. ThalamusDB (jo2025thalamusdb) explores approximate query processing for multimodal data, supporting natural language predicates over visual, audio, and textual content through zero-shot models combined with relational operators.
Several systems have proposed specialized approaches to LLM-powered SQL processing. SUQL (liu2024suql) augments SQL with answer and summary operators for knowledge-grounded conversational agents, with the focus on row-wise LLM operations for question answering. ZenDB (lin2024towards) optimizes SQL queries for extracting structured data from semi-structured documents with predictable templates. Palimpzest (liu2025palimpzest) provides a declarative framework for LLM-powered data processing with specialized map-like operations and basic cost-based optimizations. UQE (dai2024uqe) studies embedding-based approximations for LLM-powered filters and stratified sampling for aggregations, though it provides best-effort performance without accuracy guarantees.
Beyond SQL-centric approaches, several frameworks have emerged for LLM-powered data processing. Aryn (anderson2024design) offers a Spark-like API with PDF extraction and human-in-the-loop processing capabilities. DocETL (shankar2025docetl) proposes agent-driven pipeline optimization for complex document processing tasks and uses LLM agents to explore task decomposition strategies. EVAPORATE (arora2023language) specializes in extracting semi-structured tables through code synthesis. These systems demonstrate the potential of LLM-powered processing, but often lack formal cost models and provide limited performance guarantees.
Optimizing expensive predicates and user defined functions was the subject of a number of research papers in the past (hellerstein6caching) (hellestein93predicate) (chaudhuri99optimization). This line of work inspired many of the optimizations performed in AISQL. In (hellestein93predicate) a cost-based optimization technique was introduced that improved query performance by selectively reordering or “migrating” expensive selection predicates in relational query plans. The authors developed an analytical model and algorithm that balances predicate selectivity and evaluation cost, achieving faster query execution without exhaustive reordering. (hellerstein6caching) presents strategies to optimize query performance by caching the results of costly user-defined or external function calls within a query. It introduces algorithms that determine when and how to reuse previously computed method results, reducing redundant evaluations and significantly improving overall query efficiency. Finally, (chaudhuri99optimization) proposed techniques to integrate expensive, user-defined predicates into traditional cost-based query optimization frameworks. It developed a formal model for estimating predicate costs and selectivities, enabling the optimizer to reorder and combine such predicates effectively to minimize overall query execution time.
@@
8. Conclusions
The proliferation of unstructured data alongside traditional structured datasets has created a fundamental gap in modern data warehouses: users need to perform semantic reasoning over hybrid data but are forced to export data to external systems, orchestrate complex pipelines, and sacrifice the optimization benefits of integrated query processing. This paper presents Snowflake AISQL, a production SQL engine that bridges this gap by integrating LLM capabilities directly into SQL through semantic operators.
AISQL introduces six semantic operators that compose naturally with relational SQL primitives. These operators enable users to express complex analytical tasks that blend structured queries with semantic reasoning in a single declarative statement, and eliminate the need for external data movement and manual pipeline orchestration.
The core technical challenge we address is making semantic operators efficient at a production scale. LLM-based semantic operators are orders of magnitude more expensive than traditional SQL operations; their cost and selectivity are opaque to the query optimizer, and conventional optimization heuristics can produce plans that are catastrophically inefficient. Our solution comprises three techniques driven by analysis of real customer workloads. AI-aware query optimization treats LLM inference cost as a first-class objective, resulting in efficient operator placement and predicate evaluation order. Adaptive model cascades reduce costs through intelligent routing between lightweight proxy models and powerful oracle models. Query rewriting for semantic joins transforms quadratic-complexity joins into a linear multi-label classification task and, as a result, delivers speedups with improved prediction quality.
Our production deployment at Snowflake validates these design choices. Analysis of customer workloads reveals that AI operators consume the majority of execution costs, and nearly half of queries involve multi-table operations. These observations directly motivate our focus on join optimization and cost-aware planning. Extensive experimental evaluation across diverse benchmarks demonstrates that AISQL successfully delivers both efficiency and quality at the production scale.
AISQL represents a step toward unified data processing architectures where semantic reasoning over unstructured content is a first-class database capability rather than an external add-on. By integrating LLM capabilities into the SQL engine, with a focus on cost optimization, we enable organizations to unlock insights from their data assets through familiar declarative interfaces, while maintaining the performance, scalability, and reliability expected from modern data warehouses.